OpenStudy (anonymous):

subtract rad100-rad10

OpenStudy (anonymous):

\[\sqrt{10^2}-\sqrt{10}=10-\sqrt{10}\]

OpenStudy (anonymous):

Oh, that's what rad means

OpenStudy (anonymous):

rad 90 wrong answer

OpenStudy (anonymous):

Thanks for pointing that out Nancy

OpenStudy (anonymous):

urwc